A 3.5 kg model car travels down a ramp that is 2.3 m tall. At the top of the ramp, it is

moving with an initial speed of 3.4 m/s. At the bottom of the ramp is a 5,0 cm flat

surface that has a spring on the end of it, with a spring constant of 75.0 N/m.


What is the speed of the car just as it hits the spring? The ramp has a coefficieht

of friction of 0.12.


1
Expert's answer
2022-01-10T09:15:05-0500

Apply conservation of energy:


"\dfrac 12mv_i^2+mgh+\mu mgh/\sin\theta=\dfrac12mv_f^2,\\\space\\
v_f=\sqrt{v_i^2+2gh+2\mu gh/\sin\theta}."

To find the final velocity just before hitting the spring, you need to know the angle theta of the ramp above the incline.